### v2

Abstract
We investigate the generalization capabilities of small language models under two popular adaptation paradigms: few-shot prompting and supervised fine-tuning. While prompting is often favored for its parameter efficiency and flexibility, it remains unclear how robust this approach is in low-resource settings and under distributional shifts. This paper presents a comparative study of prompting and fine-tuning across task formats, prompt styles, and model scales, with a focus on their behavior in both in-distribution and out-of-distribution (OOD) settings.
Beyond accuracy, we analyze the internal representations learned by each approach to assess the stability and abstraction of task-specific features. Our findings highlight critical differences in how small models internalize and generalize knowledge under different adaptation strategies. This work offers practical guidance for model selection in low-data regimes and contributes empirical insight into the ongoing debate over prompting versus fine-tuning. Code for the experiments is available at the following link. [^1] https://anonymous.4open.science/r/moss-small-lm-experiments-BD03/moss_small_lm.ipynb
1 Introduction
Few-shot prompting and supervised fine-tuning are two widely adopted strategies for adapting pretrained language models (LMs) to downstream tasks. Prompting adapts models by conditioning on in-context examples at inference time without updating model parameters Brown et al. (2020), whereas fine-tuning involves directly optimizing the model on labeled data. While prompting is attractive for its flexibility and efficiency, its reliability in low-resource settings—particularly for small-scale language models like GPT-2 Radford et al. (2019) and DistilGPT2 Sanh et al. (2019)—remains uncertain.
In this work, we present a systematic comparison of prompting and fine-tuning using three GPT-2 variants: distilgpt2, gpt2, and gpt2-medium, evaluated across a suite of language understanding tasks. We investigate three central questions: (1) How does prompting performance scale with the number of in-context examples compared to fine-tuning under an equivalent data budget? (2) How well does each method generalize to out-of-distribution (OOD) prompt templates? (3) How stable are their internal representations across prompt variations?
We begin with prompting on synthetic multi-task tasks (sentiment, grammar correction, arithmetic, plural forms), followed by fine-tuning on IMDb sentiment classification. In addition to standard accuracy comparisons, we use t-SNE to analyze the structure of prompt and hidden-layer representations across models and prompt styles.
This study reveals important distinctions in how small LMs internalize supervision under different adaptation strategies, with implications for generalization and representational robustness in low-data regimes.
2 Scaling and Shot Count: Accuracy and Representation Geometry
To understand how model scale affects few-shot generalization, we evaluated three GPT-2 variants—distilgpt2, gpt2, and gpt2-medium—on synthetic classification tasks with varying numbers of in-context examples. Each model was prompted with $k\in\{1,2,3,4,5\}$ labeled examples and queried on a held-out test instance. The predicted label $\hat{y}$ was derived from the model’s completion of the prompt
$P=[x_{1},y_{1};\ldots;x_{k},y_{k};x_{\text{test}},\ \_?],$
where the goal is to infer the label $y_{\text{test}}$. As expected, accuracy improved with increasing $k$, with gpt2-medium consistently outperforming its smaller counterparts.
However, a t-SNE visualization  van der Maaten & Hinton (2008) of the final hidden state embeddings from each model (see Appendix 8.2, Figure 3) reveals an interesting asymmetry: although gpt2-medium achieved the highest task classification accuracy, its prompt representations did not form the most visually distinct clusters in the 2D projection. In contrast, gpt2 (base) exhibited clearer task-level separation under t-SNE, suggesting that the embeddings of larger models may become more abstract or compressed, making them less amenable to low-dimensional geometric partitioning.
This decoupling between performance and representational separability implies that scaling improves behavioral accuracy but does not necessarily enhance interpretability via dimensionality-reduced visualization. These findings highlight that few-shot generalization involves both behavioral alignment and latent structuring—two aspects that do not always co-occur in small-scale diagnostic settings.
(a) Accuracy across shot counts.
(b) Prompting vs fine-tuning.
Figure 1: Comparison of (a) few-shot accuracy trends and (b) training strategies under the same data budget.
3 Prompting vs. Fine-Tuning Accuracy
To quantify the performance gap between prompting and parameter tuning, we evaluate three GPT-2 variants on a binary sentiment classification task using the IMDb dataset. A random sample of 1000 examples is drawn and split into training and test subsets ($80\%/20\%$). Each model is exposed to a fixed budget of $k=5$ labeled examples from the training set, used under two distinct paradigms.
In the first, prompting, the examples $\{(x_{i},y_{i})\}_{i=1}^{5}$ are formatted as in-context demonstrations within a prompt. The model completes the sequence for a test query $x_{\text{test}}$ to produce a label prediction $\hat{y}$, without modifying model parameters. In the second, fine-tuning, the same examples are used to optimize the cross-entropy loss via parameter updates:
$\mathcal{L}_{\text{FT}}=-\frac{1}{k}\sum_{i=1}^{k}\log p_{\theta}(y_{i}\mid x_{i}),$
where $\theta$ denotes the model weights.
The table below summarizes accuracy scores on the held-out test set:
Model Prompting Accuracy Fine-Tuning Accuracy
distilgpt2 0.485 0.820
gpt2 0.495 0.855
gpt2-medium 0.535 0.870
Fine-tuning consistently outperforms prompting across all scales, with gains exceeding 30 absolute percentage points. This highlights the efficacy of even minimal gradient-based adaptation when data is scarce. Prompting accuracy improves modestly with model size, reflecting an increase in pre-trained capacity for in-context reasoning. However, these gains remain small relative to the improvements achieved via fine-tuning. The performance gap, defined as $\Delta=\mathcal{A}_{\text{FT}}-\mathcal{A}_{\text{Prompt}}$, is approximately constant across models and underscores that parameter tuning contributes orthogonal benefits to scaling.
These findings also carry implications for lightweight models. The smallest variant, despite having limited in-context learning ability, achieves a substantial accuracy increase through fine-tuning. This suggests that even low-capacity models retain latent flexibility when supervised adaptation is allowed. In practical scenarios such as edge deployment or privacy-sensitive fine-tuning, small models may still deliver competitive performance when gradient access is available.
In summary, while prompting offers convenience, fine-tuning remains the dominant approach in small-data regimes particularly when performance is critical and modest training is feasible.
4 Generalization to Out-of-Distribution (OOD) Prompts
We investigate how model behavior changes under variations in prompt formulation, using three semantically aligned but structurally distinct formats: standard, QA-style, and good/bad. While the underlying task—binary sentiment classification on IMDb—remains fixed, these prompt types differ in phrasing and label tokens Reynolds & McDonell (2021). Prompt construction details are provided in Appendix 8.1.
Let $x\in\mathcal{X}$, $y\in\{0,1\}$, and $\pi_{p}$ be the prompt formatter for type $p\in\mathcal{P}$. For in-context prompting, a $k$-shot input is constructed as:
$\texttt{Prompt}=\big{\|}_{i=1}^{k}\pi_{p}(x_{i},y_{i})\,\big{\|}\,\pi_{p}(x_{\text{test}},\ \text{?}),$
[…]
Prompting results, shown in Figure 2(a), reveal that all models maintain consistent accuracy between the standard and QA-style formats. However, performance drops sharply for the good/bad format, especially in gpt2-medium, suggesting strong dependence on lexical alignment Webson & Pavlick (2022). Interestingly, distilgpt2 is more stable, possibly due to reduced reliance on pretraining-specific surface cues.
We next test whether fine-tuning improves robustness to such prompt variations. Each model is fine-tuned using the standard prompt, then evaluated across all prompt types. Given a prompt $\pi_{p}(x)$, prediction is computed as:
$\hat{y}=\arg\max f_{\theta}(\texttt{Tokenizer}(\pi_{p}(x))),$
where $f_{\theta}$ is the fine-tuned model. As shown in Figure 2(b), accuracy remains high across all prompt variants. The performance gap seen in prompting vanishes, indicating that fine-tuning enables models to internalize task semantics beyond surface form, improving generalization to OOD prompts.
(a) Prompting: sensitivity to label phrasing
(b) Fine-tuning: robustness across formats
Figure 2: Accuracy across prompt formats using (a) in-context prompting and (b) fine-tuned classifiers. Fine-tuning improves robustness to surface variation.
5 Probing Internal Representations with Linear Classifiers
We investigate the extent to which different GPT-2 model variants encode task-discriminative information in their internal representations. Using a set of synthetic prompts from three task types—sentiment analysis, arithmetic addition, and factual QA—we extract final-layer hidden states from each model and evaluate whether a simple linear classifier can recover task identity.
Given an input sequence $x$, we obtain its representation $\mathbf{h}\in\mathbb{R}^{d}$ by mean-pooling the final-layer hidden states:
$\mathbf{h}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{h}_{t}^{(L)},$
where $\mathbf{h}_{t}^{(L)}$ is the hidden state of token $t$ from the final layer and $T$ is the number of tokens in the input. These representations are used to train a logistic regression classifier to predict the task label.
To assess stability, we run the experiment 10 times with different train-test splits and report the mean classification accuracy. We observe a consistent improvement with model scale: distilgpt2 achieves an average accuracy of 67%, gpt2 reaches 60%, and gpt2-medium peaks at 73%.
These results suggest that internal representations become increasingly task-aware and linearly decodable with scale. To better understand the geometric structure of these embeddings, we visualize them using t-SNE in the Appendix (Figure 4). The plots confirm that larger models form more compact and separable clusters for each task category, providing qualitative support for the quantitative probe accuracy results.
6 Layer-wise Probing of Task Representations
To understand how task-discriminative information emerges across the model depth, we perform layer-wise probing on each GPT-2 variant. For each transformer layer, we extract the mean-pooled hidden state from synthetic task inputs and train a logistic regression classifier to predict the task category (sentiment, arithmetic, QA). This allows us to evaluate the linear decodability of task identity at each layer.
Figure 5 shows the average probe accuracy per layer, aggregated over 10 runs. We observe that deeper layers consistently encode more linearly separable task information. For gpt2-medium, accuracy climbs sharply after the midpoint and saturates in the final layers. In contrast, distilgpt2 achieves peak probe accuracy in the middle layers, suggesting a compressed representational depth.
These results support the view that representational abstraction accumulates with depth and scale, and that later layers consolidate semantic structure in a form more accessible to simple classifiers.
7 Conclusion and Future Directions
This study demonstrates that small-scale diagnostic experiments can reveal nuanced trends in language model representations and generalization. We show that task identity becomes more linearly decodable with scale and depth, yet visualization via t-SNE reveals that larger models do not always exhibit cleaner geometric separation—highlighting a disconnect between abstraction and interpretability. Our probing results confirm that later layers encode more task-relevant information, and fine-tuning yields greater robustness to prompt variation than in-context learning alone.
Future work may explore richer probing techniques (e.g., CKA, RSA), test robustness under syntactic or cross-lingual shifts, and investigate whether modular or retrieval-augmented models can improve both interpretability and performance at smaller scales. These directions support the case for using small, controlled setups to build principled understanding of model behavior.